### v1

# Magneto: Combining Small and Large Language Models for Schema Matching

Abstract.
Recent advances in language models opened new opportunities to address complex schema matching tasks. Schema matching approaches have been proposed that demonstrate the usefulness of language models, but they have also uncovered important limitations: Small language models (SLMs) require training data (which can be both expensive and challenging to obtain), and large language models (LLMs) often incur high computational costs and must deal with constraints imposed by context windows. We present Magneto, a cost-effective and accurate solution for schema matching that combines the advantages of SLMs and LLMs to address their limitations. By structuring the schema matching pipeline in two phases, retrieval and reranking, Magneto can use computationally efficient SLM-based strategies to derive candidate matches which can then be reranked by LLMs, thus making it possible to reduce runtime without compromising matching accuracy. We propose a self-supervised approach to fine-tune SLMs which uses LLMs to generate syntactically diverse training data, and prompting strategies that are effective for reranking. We also introduce a new benchmark, developed in collaboration with domain experts, which includes real biomedical datasets and presents new challenges to schema matching methods. Through a detailed experimental evaluation, using both our new and existing benchmarks, we show that Magneto is scalable and attains high accuracy for datasets from different domains.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/VIDA-NYU/data-harmonization-benchmark and https://github.com/VIDA-NYU/data-integration-eval.
1. Introduction
The rapid increase in the volume of structured data– from data published in scientific articles (Nature, 2024; Science, 2024) and repositories (National Cancer Institute, 2024a; pdc, 2023; ukbio, 2024; European Organization For Nuclear Research and OpenAIRE, [n.d.]) to open government portals (of New York, 2024; datagov, 2024) – creates new opportunities to answer important questions through analytics and predictive modeling. But often, data from multiple sources must be integrated to answer these questions. Consider the following example.
Example 1.1.
Proteogenomic analysis of endometrial cancer tumors: In the context of the National Cancer Institute’s Clinical Proteomic Tumor Analysis Consortium (CPTAC) (cptac, 2024), Li et al. (2023) carried out a comprehensive proteogenomic analysis of endometrial cancer tumors. They pooled data from ten studies (available as supplementary material in the original primary research papers) that cover multiple patient cohorts and cancer types. To facilitate the integration and analysis, they mapped each dataset into the GDC standard, a data standard set by the National Cancer Institute’s Genomic Data Commons (GDC) for cancer genomic data (Heath et al., 2021; National Cancer Institute, 2024a). Even though the studies had been carried out by members of the CPTAC effort, the datasets containing patient case and sample data used many different representations for variable names and values. Integrating these data required a substantial effort to match variables from each dataset source schemata to the target GDC format, which encompasses over 700 attributes.
Even though the schema matching (SM) problem has been extensively studied in the data management literature (Doan et al., 2012; Cafarella et al., 2009; Miller, 2018; Koutras et al., 2021), we lack solutions that are general and work well for data in the wild. Schema matching still requires a time-consuming, manual curation process for complex matching tasks like the one described above (Adhikari et al., 2021; Cheng et al., 2024). Matching is particularly challenging in the presence of ambiguity and heterogeneity in the representation of attributes and values (see Table 1). SM approaches that rely on attribute names, data types and values for similarity assessments are likely to fail for such matches. As a point of reference, we assessed the effectiveness of state-of-the-art schema matching strategies for the biomedical datasets from Example 1.1. As shown in Figure 1, they perform poorly: the best technique achieves at most 0.45 of mean reciprocal rank (MRR) and requires powerful computational resources to run. We discuss these results in detail in Section 6.
Figure 1. Trade-off between runtime and accuracy measured as the mean reciprocal ranking (MRR) for different schema matchers for the matching task of Example 1.1. Magneto variants (in orange) achieve higher MRR and better accuracy-runtime trade-offs than traditional methods (in blue).
Determining columns correspondences may require knowledge beyond the schema and contents of a table. Table 1 shows some possible matches for the proteogenomic analysis. Sometimes, selecting the correct match is difficult even for subject matter experts. For example, for the attribute patient_age from one of the datasets, there are at least three plausible matches in GDC: age_at_diagnosis, days_to_birth, and age_at_index. Without additional context, it is difficult to determine the correct match.
Schema Matching and Language Models. There has been renewed interest in data integration given the new capabilities introduced by language models (Feuer et al., 2024; Narayan et al., 2022). For schema matching, approaches have been proposed that have shown promising results (Castro Fernandez et al., 2018; Cappuzzo et al., 2020; Du et al., 2024; Tu et al., 2023). A key challenge in schema matching is the estimation similarities between two columns. Pre-trained Language Models (PLMs) ,referred to as SLMs in this paper to distinguish them from large language models (LLMs), support the creation of representations, or embeddings, for columns enriched with semantic information. The similarity between two embeddings can serve as a proxy for column-matching scores (Castro Fernandez et al., 2018; Cappuzzo et al., 2020). SLMs have also been fine-tuned for schema matching (Du et al., 2024) and general matching tasks (Tu et al., 2023), leading to accuracy gains compared to the general models.
LLMs are trained using large, generic data corpora and thus contain knowledge that can assist in obtaining additional semantics necessary to identify matches. Prompting strategies combined with fine-tuned models have been shown effective to improve table understanding and help in integration tasks (Li et al., 2024a).
While prior work has shown the usefulness of SLMs and LLMs for SM, they also present significant practical practical challenges:
  • $\bullet$
    Challenge 1: Fine-tuning SLMs can lead to significant performance improvements, but this requires the availability of manually curated training data, which can be expensive to create.
  • $\bullet$
    Challenge 2: LLMs usually do not need fine-tuning, but they have a fixed context window: if a prompt exceeds this limit, it may need to be truncated before it is sent to the model. Since small context windows are common in LLMs, this can lead to losing critical context and limit their use for large schemata and tables. While recent LLMs are starting to support larger windows, their cost scales with input and output size. Besides, recent studies have shown that the accuracy of LLMs can significantly decrease when given long prompts for complex tasks (Li et al., 2024b). Last but not least, calls to LLM APIs can have high latency, especially for large prompts.
Table 1. Column names and values for candidate matches between columns in source tables and the GDC format. These illustrate challenges involved in identifying matches in the presence of heterogeneity in the schemata. Note: GDC does not specify values for numerical attributes such as age.
Source Column and Values Target Candidate Columns
Histologic_Grade_FIGO: FIGO grade 1, FIGO grade 2, FIGO grade 3 tumor_grade: G1, G2, G3
who_nte_grade: G1, G2, G3
adverse_event_grade: Grade 1, Grade 2, Grade 3
Path_Stage_Primary_Tumor-pT: pT1a (FIGO IA), pT2 (FIGO II), pT3b (FIGO IIIB) uicc_pathologic_t: T1a, T2, T3b
ajcc_pathologic_t: T1a, T2, T3b
Path_Stage_Reg_Lymph_Nodes-pN: pN1 (FIGO IIIC1), pN0, pNX ajcc_pathologic_n: N1, N0, NX
uicc_pathologic_n: N1, N0, NX
Age: 65, 72, 49 age_at_diagnosis: n/a
days_to_birth: n/a
age_at_index: n/a
Our Approach. We introduce Magneto, a framework that combines small (pre-trained) language models (SLMs) and large language models (LLMs) to derive cost-effective and general solutions for SM. As illustrated in Figure 2, Magneto is structured in two phases: candidate retrieval selects a subset of the possible matches; and reranker, which ranks the candidates to make it easier for users to examine and select matches. We use this framework to explore three new strategies that use SLMs and LLMs: Using SLM as the candidate retriever and LLMs as the reranker; an automated, LLM-powered fine-tuning process for SLMs; and a combination of these. We propose new methods that address the context window and cost challenges associated with LLMs and the creation of training data to fine-tune SLMs.
To address Challenge 1, we leverage LLMs to automate the process of fine-tuning SLMs. Instead of relying on manually created training data and structure-based augmentation (e.g., row shuffling and sampling (Fan et al., 2023; Du et al., 2024)), we use LLMs to derive data. Using LLMs, we can add syntactic diversity and capture different representations for column names and their values.
We use a multi-pronged approach to address Challenge 2. By structuring the matching process in two phases, Magneto can use cheaper SLM-based methods for filtering candidates, thus reducing the number of matches that need to be checked by more costly rerankers, e.g., LLM-based rerankers, thereby reducing the cost and latency. We must also ensure that all necessary details are included in the prompt while staying within the context window limit. Thus we need to create representative samples the contents of the columns in candidate matches to be ranked by the LLM. In addition, LLMs are designed for textual data. To use them for tabular data, we must serialize the column content. Selecting the right serialization strategy is still an open research problem that has attracted substantial attention (Fang et al., 2024; Dong and Wang, 2024). We explore different alternatives for both sampling and serialization.
Our ultimate goal is to devise strategies that have high accuracy and low cost. However, there is often a trade-off: high-accuracy is associated with higher costs and running times. Figure 1 illustrates the trade-offs between runtime and accuracy for existing schema matching algorithms and Magneto. Traditional methods tend to cluster at relatively high accuracy with long runtimes or relatively low runtimes with reduced accuracy. In contrast, Magneto variants strike a balance across both dimensions and have the advantage of requiring no prior training data.
We conduct an extensive experimental study of Magneto on various datasets from two benchmarks. The first is a new benchmark, that we developed in collaboration with domain experts, which includes real biomedical datasets. Our experiments show that many existing SM solutions struggle with this benchmark, particularly given the complexity and syntactic variability present in its datasets. In contrast, Magneto demonstrates superior accuracy and generally faster runtimes. We also evaluate Magneto on the Valentine benchmark (Koutras et al., 2021) and observe that it also performs well for the Valentine datasets which cover a variety of domains.
Figure 2. Magneto overview. Given source and target tables, the framework identifies matches in two phases: A small language model, automatically fine-tuned using LLM-generated data, retrieves and ranks candidate matches, which are then assessed and reranked by an LLM. This process makes it possible to efficiently identify matches and generalize across domains and datasets. Different schema matching strategies can be derived by adopting different combinations of methods for these phases.
Contributions. Our main contributions are summarized as follows:
  • $\bullet$
    We introduce the Magneto framework that effectively combines SLMs and LLMs and allows the creation of SM strategies that attain a balance between accuracy and runtime for diverse data and SM tasks (Section 2).
  • $\bullet$
    We propose an LLM-powered method that automatically generates training data needed to fine-tune SLMs, making it possible to effectively use these low-cost models for specific SM tasks. We also present a fine-tuning pipeline that uses contrastive learning and, by combining triplet loss and online triplet mining, effectively utilizes the generated data to enhance the discriminative power of column embeddings for SM (Section 3).
  • $\bullet$
    Unlike previous approaches that use LLMs for SM, we show that reranking candidate matches with LLMs, derived from less costly methods, achieves high accuracy at lower costs. (Section 4).
  • $\bullet$
    In collaboration with biomedical researchers, we created a new benchmark that represents a real schema matching effort (Li et al., 2023), encompassing characteristics not present in existing benchmarks and presents new challenges for SM methods (Section 5).
  • $\bullet$
    We perform an extensive experimental evaluation of different strategies derived using our framework, comparing them against SM methods over datasets that cover a wide range of domains. We also carry out ablation studies to assess the effectiveness of different choices for column sampling and serialization. (Section 6).
2. Motivation and Solution Overview
Biomedical Data Integration: Opportunities and Challenges. There have been substantial investments in infrastructure to share biomedical data (National Institutes of Health, 2020). Today, many datasets are available in repositories and published with scientific articles (pdc, 2023; National Cancer Institute, 2024a). The ability to pool these data can advance research in multiple diseases and across different populations. However, established technologies are limited in their ability to integrate these data. Our research was motivated by this challenge as part of a collaboration with biomedical researchers in the context of the ARPA-H Biomedical Data Fabric (BDF) program (ARPA-H, 2024), which aims to develop data integration and usability technologies.
We conducted formative interviews with biomedical experts and found that they rely heavily on manual processes and dataset-specific scripts to perform schema matching (Mertins et al., 2016; Li et al., 2023; Cheng et al., 2024). The example described in Section 1 illustrates some of the difficulties encountered in one study. Besides the time required, the process they follow is error-prone and difficult to reproduce. As data integration is a key step in most of their studies, they would benefit from automated (and semi-automated) techniques that streamline integration and are also transparent and reproducible.
We have identified two important requirements based on the researchers’ needs. They need approaches that: generalize across different kinds of data (e.g., genomics, proteomics, electronic health records), for different diseases (e.g., different types of cancers, auto-immune and rare diseases), from many sources (e.g., publications, hospitals, data commons); and that facilitate curation, since in many cases it’s difficult (even for experts) to determine correct matches. These requirements have guided the design of Magneto.
2.1. Definitions and Evaluation Metrics
Before we describe our approach, we introduce the terminology used in the remainder of this paper. We formally define the problem we are addressing and the performance metrics we use to evaluate the methods we propose for schema matching.
Definition 2.0.
(Schema Matching) Let $\mathbf{S}(A_{1},\dots,A_{n})$ be a source table and $\mathbf{T}(B_{1},\dots,B_{m})$ be a target table, where $A_{i}\in\mathbf{S}$ and $B_{j}\in\mathbf{T}$ are columns that define the schemata. Each column $A\in\mathbf{S}\cup\mathbf{T}$ has an associated domain, denoted $\mathcal{D}(A)$, representing the set of possible values that the column can take; note that $\mathcal{D}(A)$ may be empty. Schema matching focuses on aligning the table schemata by establishing correspondences between columns representing the same real-world concept or entity. A matching algorithm (or matcher) aims to identify pairs $(A_{i},B_{j})$ that represent the same (or semantically equivalent) column based on various factors, such as their domains and names. Thus, a matcher $\mathcal{M}$ can be seen as a function that generates a schema mapping $M\subseteq\mathbf{S}\times\mathbf{T}$, where each element $(A,B)\in M$ represents a correspondence between a source column $A$ and a target column $B$, where $\mathcal{D}(A)\approx\mathcal{D}(B)$, meaning that the domains of columns $A$ and $B$ are related or overlap. ∎
Matching algorithms often associate a score with each match they derive. These scores are used to generate ranked lists containing the derived matches, which help users explore matches by prioritizing the highest-scoring candidates. These lists provide a global ranking of best matches $(A_{i},B_{j})$ among all pairs of possible matches of a given pair of tables S and T or a per-column ranking of the best matches $B_{j}$ for a given source attribute $A_{i}\in\textbf{S}$. Thus, a common approach to evaluate schema matching methods is to assess their ability to produce high-quality ranked lists of matches (Koutras et al., 2021; Du et al., 2024). We use two measures in our evaluation, Mean Reciprocal Rank (MRR) and Recall at Ground-Truth Size (Recall@GT), which capture different aspects of matching methods: user experience and result completeness.
Definition 2.0.
(Mean Reciprocal Rank) Let $matches[A]$ denote the ranked list of matches produced by a schema-matching algorithm for the source column $A\in\textbf{S}$, and $r_{A}$ be the position of the first correct target column $B$ within the ordered list $matches[A]$. The reciprocal rank (RR) of an individual column $A$ is the multiplicative inverse of the rank, i.e., $\frac{1}{r_{A}}$. For instance, the RR is $1$ if the match is found in the first place, $\frac{1}{2}$ for the second place, $\frac{1}{3}$ for the third place, and so on. The mean reciprocal rank (MRR) for a table S is the average RR over the subset of columns $\textbf{S}^{\prime}$ that contain a correct match in the ground truth:
(1) $\text{MRR}=\frac{1}{|\textbf{S}^{\prime}|}\sum_{A\in\mathbf{S}^{\prime}}\frac{1}{r_{A}}.$
Intuitively, MMR measures how long it takes to find the first relevant match when examining the ranked list of matches. ∎
MRR is commonly used to evaluate methods that produce a ranked list of items, including the evaluation of search engine results and responses derived by question-answering systems (Schütze et al., 2008; Voorhees et al., 1999). In short, MRR evaluates the quality of the ranking. The prioritization of the top-ranked correct result aligns with an enhanced user experience. In our problem, this corresponds to ranking the most relevant match for a given source column as highly as possible: a high MRR score indicates that a user can find the correct match more easily when judging multiple matches for a column.
We also use the recall at ground truth size (Recall@GT), a standard measure used in recent schema matching works (Koutras et al., 2021). In contrast to MRR, Recall@GT is computed over a global ranking of the matches, i.e., instead of considering the ranking for each source column, it considers a merged list of all matches produced.
Definition 2.0.
(Recall@GT) Let $matches$ denote the global ranked list containing all matches produced by a matcher that considers all pairs of possible matches between columns from S and T, and let $\mathcal{M}$ denote a set containing only the top-$k$ best results in $matches$. Moreover, let $\mathcal{G}$ be the set of ground truth matching pairs $(A,B)$. Recall@GT measures the fraction of relevant matches in the ground truth that also appears in $\mathcal{M}$, where $k$ is given by the size of the ground truth, $k=|\mathcal{G}|$. More formally,
[…]
Intuitively, it measures how well the matcher can place all correct matches of a table S at the top of the global ranked list. ∎
2.2. Magneto: Overview
Magneto consists of two main components, a candidate retriever and a reranker, which use small and large language models to efficiently derive accurate matches. The intuition is to first apply a cheaper approach to retrieve and filter candidates so that a more sophisticated (and expensive) methods can be used to accurately identify the correct matches from a smaller candidate set. By doing so, we reduce the overall cost of schema matching and, at the same time, attain high accuracy. Figure 2 shows one variant of Magneto that uses a fine-tuned language model as the retriever, and a large language model as the reranker. The use of language models enables Magneto to serve as general solution that supports schema matching across multiple domains.
Candidate Retriever. The candidate retriever leverages SLMs to generate a ranked list of potential matches from the target table for each input column. It uses column embeddings to estimate column pair similarities (Cappuzzo et al., 2020; Fan et al., 2023). SLMs are a good choice for this step given their ability to capture semantic similarity and efficiency.
Pre-trained (generic) SLMs such as BERT (Devlin, 2018), RoBERTa (Liu, 2019), and MPNet (Song et al., 2020), can struggle with nuanced syntactic differences and lack contextual knowledge for complex schema matching tasks involving domains not represented in their training data. To perform complex tasks effectively, these models need to be fine tuned. We propose to leverage LLMs to automatically generate the data needed to fine tune SLMs. We show that LLMs can help derive high-quality training data that reflect instances of variability for semantically similar columns that arise in real data. This approach, which we describe in Section 3, improves the robustness of the SLM-based retriever, making it possible for it handle complex matches without requiring human-labeled training data. We refer to Magneto configurations that use this approach using the label ft. Note that Magneto-ft only invokes LLMs once, during the training phase for a given domain. Once trained, the fine-tuned SLM can perform multiple inferences efficiently without making calls to LLMs.
Match Reranker. While fine-tuning improves the performance of SLMs, it is not sufficient to handle schema matching tasks that include domains and heterogeneity not covered during the fine tuning. To improve generalizability, Magneto uses LLMs as rerankers to refine the candidates identified by the SLM-based retriever. This approach, which we refer to as Magneto-llm, enhances accuracy by leveraging carefully designed prompts and techniques that enable the LLM to judiciously assess matches and discern subtle semantic nuances, which are challenging for the SLM to detect independently. As discussed in Section 4, Magneto-llm also strives to minimize costs by optimizing LLM resources.
Varying Retrievers and Rerankers. The architecture of Magneto allows the combination of different alternatives for retrievers and rerankers. This is useful for studying the tradeoffs of various approaches and to derive solutions that match the accuracy requirements and computational constraints for different applications.
For the candidate retriever, we can combine traditional matching techniques with the embedding strategies. For example, when columns in a match have the same name, module minor variations in case and punctuation, we can assign a perfect similarity score of 1.0 to the match. During our experiments, this simple technique consistently improved overall accuracy, albeit slightly, in most cases.
We implemented an alternative match reranker, the bipartite-graph reranker, which adapts the a filtering technique from (Melnik et al., 2002). This approach offers an algorithmic alternative for reranking, making it particularly suitable for scenarios where LLMs are unavailable or where strict runtime constraints must be met. This solution combines all ranked match candidates of each source column into one global ranked list and transforms it into an undirected bipartite graph where the nodes represent columns. There are two (disjoint) node sets consisting of the columns in the source and target tables (respectively). A match is represented as an edge in the graph, with a corresponding score as its edge weight. To identify the best match for the set of attributes in the source column, it uses the algorithm from (Crouse, 2016), which solves the assignment problem in polynomial time and is efficient even for large graphs. We prioritize the matches in the assignment by placing the corresponding edges at the top of the list. To account for the remaining matches, we move them to the bottom of the list while preserving their relative order..
3. Using LLMs to Fine-Tune SLMs
SLMs for Candidate Match Retrieval. Pre-trained small language models (SLMs) have been used to help with schema-matching related tasks by encoding semantic information from column names and values into dense vector representations–embeddings (Tu et al., 2023; Zhang et al., 2023; Fan et al., 2023; Cong et al., 2023; Dong et al., 2023; Du et al., 2024). To identify matches, embeddings of source and target columns can be compared using cosine similarity: high similarity scores indicate a higher likelihood that the columns match.
SLMs are pre-trained on and work well for general natural language tasks, but their ability to interpret tabular data is limited. These models must often be fine-tuned to handle tasks that involve tables. However, fine-tuning approaches require large amounts of labeled data for training (Tu et al., 2023; Dong et al., 2023). Unfortunately, this is impractical in many scenarios such as the integration of biomedical data (Example 1.1), for which training data is hard to obtain.
In the absence of training data, it is possible to apply augmentation techniques to automatically generate variations of data to be used as positive examples. For example, given a column, different versions can be generated by shuffling rows, sampling values, and applying perturbations to values (Fan et al., 2023; Dong et al., 2023). However, these variations may not fully capture the heterogeneity found in real data.
We introduce a new method that leverages LLMs to generate training data and present a pipeline to fine tune SLMs for schema matching (Section 3.2). Another important consideration in the use of SLMs, that we discuss this in Section 3.1, is how to represent tables and their columns.
Figure 3. LLM-Powered Fine-tuning Pipeline
3.1. Column Serialization and Value Sampling
Column Serialization. Since SLMs interpret inputs as sequences of tokens (i.e., regular text), we must transform each column into a token sequence that the model can process. Recent research has explored various approaches to serializing columns into a textual format (Badaro et al., 2023). Here, we explore the impact of different approaches to column serialization for schema matching.
Given column $A$ with column type $type(A)$ and sample values $v_{1},v_{2},\dots,v_{m}$, we consider the following approaches:
[…]
$\mathcal{S}_{\texttt{verbose}}(A)$ $=\texttt{[CLS]}\text{Column: }A\,\texttt{[SEP]}\,\text{Type: }type(A)$
[…]
where the [CLS] token is used as a special symbol to indicate the start of a column, and [SEP] separates the components within a serialized column.
$\mathcal{S}_{\texttt{default}}$ is a variant of a widely-adopted serialization strategy (Suhara et al., 2022). Instead of employing commas to separate sample values, we utilize the [SEP] token. This modification prevents the model from interpreting the values as an ordered text sequence, thereby treating them as unordered, discrete features. This change is crucial in schema matching as it emphasizes the structured and independent nature of the data in each column, enabling more accurate alignment of similar but non-identical schemata.
$\mathcal{S}_{\texttt{verbose}}$ is an extension of $\mathcal{S}_{\texttt{default}}$ where prefixes are added to delineate each component (i.e., column name and values) and provide additional context for the SLM.
$\mathcal{S}_{\texttt{repeat}}$ repeats the column name multiple times to reinforce its importance, nudging the model to prioritize column names. It attempts to mitigate attention drift in SLMs, which can disproportionately focus on values rather than on column names, especially in zero-shot settings. Specifically, we chose $k=5$ because it strikes a balance between reinforcing column names without overwhelming the model with redundancy.
In addition to column name and values, we incorporate data types into the column representation. For column type inference, we classify columns into basic types: numerical, categorical, date, or binary. We classify columns with a high proportion of unique values (e.g., over 90% distinct values) as “key” columns, as they often represent unique identifiers. When column value is unavailable or type detection fails, we label the column type as “unknown”.
Priority Sampling for Column Values. To efficiently retrieve candidate matches, we generate embeddings by including a sample of column values in the serialized column representation. Different sampling techniques have been explored for various table management tasks, such as column type annotation (Feuer et al., 2024) and table union search (Fan et al., 2023). A common approach is to use weighted sampling and assign higher weights to more frequent values. We adopt this approach and incorporate coordination into the sampling process with priority sampling (Daliri et al., 2024). Priority sampling was used for inner product sketching and aimed to maximize the likelihood of selecting corresponding values across vectors, emphasizing those with larger magnitudes.
In our setting, priority sampling is adapted to increase the probability of sampling similar values during the fine-tuning and inference stages. Specifically, we use a random seed $s$ to select a uniformly random hash function $h:\{1,\dots,N\}\to[0,1]$, where $N$ represents the maximum number of unique values across all columns. For each value $v_{i}$, we compute a rank:
$R_{i}=\frac{\text{freq}(v_{i})}{h(v_{i})},$
where $\text{freq}(v_{i})$ denotes the frequency of $v_{i}$. Then, we select the sample’s first $m$ values with the largest priorities $R_{i}$.
In Magneto, we use serialization as a hyperparameter. We study the impact of different serialization strategies for schema matching in Section 6, where we also investigate different approaches to sampling column values. We experimented with varying column value sample sizes (from 10 to 30) and observed only marginal differences. Interestingly, smaller sample sizes (e.g., ten values) sometimes outperformed larger ones, likely because of reduced noise. Thus, we fixed the sample size to 10 as a default.
3.2. LLM-Powered Fine Tuning
By fine-tuning a model, our goal is for it to learn column representations so that its embeddings can spatially cluster related columns. This spatial arrangement helps the retrieval of semantically related representations using cosine similarity. With this goal in mind, the basic intuition behind our training methodology is to minimize the distance between embeddings of identical or semantically-related columns while maximizing the distance between distinct columns.
LLM-derived Training Data. As discussed above, an important challenge in fine-tuning lies in obtaining high-quality training data. With the goal of capturing syntactic heterogeneity common in real datasets, we introduce LLM-based augmentation (llm-aug), an approach that generates variations of columns that are semantically equivalent but syntactically different. The derived columns serve as the training data fine-tuning process, which adopts a self-supervised contrastive learning approach (Figure 3).
Synthetic columns are derived from an input (anchor) column as follows: 1) The anchor name and a sample of its values to the LLM using the structured prompt depicted in Figure 4, and 2) the LLM outputs columns that are semantically similar to the anchor. Examples of synthetic columns are shown in Table 2. This method establishes a class of columns considered matches (derived from the same anchor); columns within the same class serve as positive examples and those from different classes as negative examples.
Instruction Given the table column [Column Name] with values [Sample Column Values], generate three alternative column names that adhere to typical database naming conventions such as underscores and abbreviations. Additionally, provide distinct, technically correct synonyms, variants, or abbreviations for the listed values. For columns with numerical or datetime data, generate random numbers or dates appropriate to the column’s semantic meaning.
Format Ensure that each set does not exceed 15 values. Format your output as follows: alternative_name_1, value1, value2, value3, …; alternative_name_2, value1, value2, value3, …; alternative_name_3, value1, value2, value3, … Ensure your response excludes additional information and quotations.
Figure 4. Training Data Generation Prompt.
Table 2. Examples of the provided data to LLM-based reranker and the generated data.
Original Data Generated Data
Column: tissue_source_sites tumor_site
Values: Thyroid, Ovary Thyroidal, Ovarian
Column: exon gene_segments
Values: exon11, exon15 segment11, segment15
Column: masked_somatic_mutations genetic_variants
Values: MET_D1010N, FLT3_ITD D1010N_MET, ITD_FLT3
Column: max_tumor_bulk_site primary_tumor_location
Values: Maxilla, Splenic lymph nodes Maxillary, Splenic_nodes
We combine the LLM-based augmentation with other structure-based augmentation augmentation methods (struct-aug) such as random sampling and shuffling of values, and minor perturbations to the column name, including random character replacements or deletions (Fan et al., 2023). These techniques inject variability while preserving syntactic and structural similarity to the original column. By using both augmentation strategies, the model can learn to identify matches with different characteristics.
Note that we restrict this fine-tuning and training process to the target table columns to avoid inaccuracies. Applying this process to source tables with unknown true matches to target tables could mistakenly classify them as negative examples, introducing errors into the embedding space.
Triplet Loss and Online Triplet Mining. To leverage the generated synthetic columns, we implement a contrastive learning framework using triplet loss and online triplet mining (Schroff et al., 2015). Triplet loss improves upon previous contrastive loss functions by more directly optimizing for relative similarities, which, as we discovered in our experiments, yields more precise schema alignments.
In the training pipeline, each triplet consists of an anchor column $a$, a positive column $p$, and a negative column $n$; $a$ and $p$ are semantically related (i.e., they belong to the same class), while $n$ is drawn from a different class. The goal is to learn an embedding space such that the distance between the $a$ and $p$ is minimized while the distance between $a$ and $n$ is maximized. Specifically, we incorporate Batch Hard triplet loss (Schroff et al., 2015) to refine the learning of discriminative embeddings for column classification. Given $P$ classes and $K$ columns per class, the loss function is defined as follows:
(3) $\mathcal{L}=\sum_{i=1}^{P}\sum_{a=1}^{K}\big{[}\max\big{(}0,\,m+\max_{p=1\dots K}D(a,p)-\min_{\begin{subarray}{c}j=1\dots P\\ n=1\dots K\\ j\neq i\end{subarray}}D(a,n)\big{)}\big{]}$
Here, $a$ is the anchor column embedding from a specific class $i$, $p$ is the positive column embedding within the same class as the anchor, and $n$ is the negative column embedding from a different class. The function $D$ denotes the distance metric, cosine similarity in our case, which measures the similarity between embeddings. The margin $m$ is a hyperparameter that ensures a minimum distance between the positive and negative embeddings relative to the anchor, thereby enhancing the separability between classes. This loss function iterates over all classes ($P$) and multiple examples per class ($K$) within each training batch, optimizing the embeddings to emphasize inter-class distinctions and intra-class similarities.
Additionally, we apply Online Triplet Mining (Schroff et al., 2015; Hermans et al., 2017), which enhances the learning process by dynamically selecting the most challenging positive and negative examples within each training batch. This technique prioritizes triplets that maximize learning efficiency, specifically focusing on:
  • $\bullet$
    Hardest Positives: Triplets where the distance between the anchor and the positive is maximized: $D(a,p)$ is maximized such that $D(a,p)<D(a,n)$
  • $\bullet$
    Hardest Negatives: Triplets where the distance between the anchor and the negative is minimized, yet still greater than the anchor-positive distance plus the margin: $D(a,n)$ is minimized but still $D(a,n)>D(a,p)+m$
  • $\bullet$
    Semi-Hard Negatives: These are negatives that are not the closest to the anchor but are still closer than the positive plus the margin: $D(a,p)<D(a,n)<D(a,p)+m$
These conditions help the model not settle for easy examples and instead learn to distinguish subtle differences, developing more robust and discriminative features to distinguish columns. By combining triplet loss and online triplet mining, our model can effectively utilize the synthetic columns to enhance the discriminative power of the embeddings for schema matching.
Model Selection for Schema Matching. To select the optimal fine-tuned model during training, we must define an effective validation metric specifically tailored for effective schema matching tasks. Our selection process uses the measures described in Section 2, mean reciprocal rank (MRR) and recall at ground truth (Recall@GT), to evaluate model performance on synthetic data. Since no ground-truth data is available, we rely on synthetic datasets to simulate real-world scenarios. These metrics are specifically chosen to ensure that the model identifies correct matches and ranks them in a manner that reflects their true relevance.
The intuition behind using these metrics, rather than a traditional loss metric, lies in their ability to reflect the practical requirements of schema matching. Loss metrics typically provide a general sense of model error but do not differentiate between the types of errors that are more consequential for the task. In schema matching, failing to rank a highly relevant match at the top (as evaluated by MRR) or missing a relevant match entirely (as evaluated by recall) can be more detrimental than the errors captured by average loss.
[…]
4. LLMs as Rerankers
SLMs can serve as efficient retrievers, but they may fail to capture complex semantic relationships. Moreover, models fine-tuned with LLM-derived data are inherently domain-specific. To mitigate these limitations, Magneto incorporates a Large Language Model (LLM) as a reranker. Unlike existing methods that adjust match rankings based on heuristic rules and similarity metrics scores (Madhavan et al., 2001; Melnik et al., 2002; Do and Rahm, 2002; Gal, 2006; Radwan et al., 2009; Alserafi et al., 2020), Magneto leverages the knowledge accumulated by the LLM to complement the initial retrieval and improve the overall performance of SM.
A natural question arises: Why not just use an LLM for SM? While there are approaches that employ LLMs for SM (Li et al., 2024a; Parciak et al., 2024), their applicability in scenarios that involve many or large tables is limited due to high computational costs and challenges related to context windows (Challenge 2, Section 1). We posit that LLMs must be used judiciously and designed Magneto accordingly.
The core of our reranking approach is a carefully designed prompt template that converts the abstract task of column-pair similarity assessment into a more structured and interpretable process. We designed this prompt inspired by recent lessons learned from column type annotation (Feuer et al., 2024) and trends in prompt engineering (White et al., 2023; Narayan et al., 2022). Figure 5 shows the structure of the prompt.
Scoring-Oriented Instruction From a score of 0.00 to 1.00, judge the similarity of the candidate column in the source table to each target column in the target table. Each column is represented by its name and a sample of its respective values, if available.
One-shot
Example
Example:
Candidate Column:
Column: EmpID, Sample values: [100, 101, 102]
Target Columns:
Column: WorkerID, Sample values: [100, 101, 102]
Column: EmpCode, Sample values: [00A, 00B, 00C]
Column: StaffName, Sample values: [“Alice”, “Bob”, “Charlie”]
Response: WorkerID(0.95); EmpCode(0.30); StaffName(0.05)
Format Provide the name of each target column followed by its similarity score in parentheses, formatted to two decimals, and separated by semicolons. Rank the column-score pairs in descending order. Exclude additional information and quotations.
Input Source Column:[Serialized Source Column]
Target Columns:[Serialized Candidate Columns]
Response:
Figure 5. Schema Matching Prompt.
Scoring-Oriented Prompt Design. The Scoring-Oriented Instruction is a critical component of the prompt that aligns the model with the schema matching task. Specifically, this instruction involves sending one source column and the top $k$ candidate target columns retrieved by the SLM to the LLM. Previous works, such as (Li et al., 2024a) and (Parciak et al., 2024), typically employ table-wise prompts that ask LLMs to generate a ranked list without providing individual scores, which poses scalability challenges when entire tables are involved. Furthermore, column-wise ranking without scoring prevents effective table-wise comparisons, such as evaluating the recall of matches against the ground truth (recall@GT) for each pair of source and target columns.
To address these limitations, we propose a novel prompt design that requires the model to assign scores from 0.00 to 1.00 to each column pair, rather than merely producing a ranked list. This not only facilitates direct comparisons across different pairs but also allows the model to adopt a holistic view by considering all $k$ target columns simultaneously during scoring. Additionally, our serialization of the column data clarifies to the model that a text sequence comprising a column name followed by its values represents a column in a relational table. This assists in more accurately scoring matches based not only on the semantic and contextual relevance but also on the structural characteristics of the column data.
Few-Shot In-Context Learning. By providing a few examples of a task to an LLM, few-shot learning can lead to significant improvement in LLM performance (Brown, 2020; Dong et al., 2022). This method has been shown effective in various data management tasks, enabling robust performance without the need for extensive fine-tuning (Hegselmann et al., 2023; Shraga et al., 2020).
In the One-shot Example strategy, we provide a single example of the schema matching task, which outputs a ranked list with scores. This approach clarifies both the objective of the task and expected output format, as well as establishes a uniform scoring standard, enhancing the comparability of matches across all source and target column pairs. Utilizing only one example helps maintain a lightweight prompt structure, crucial for minimizing input and output token counts. The number of tokens directly impacts runtime and cost, and it can also influence accuracy (Li et al., 2024b).
Optimization of Model Cost. For the implementation of Magneto, we use GPT-4o-mini, a low-cost LLM, which has shown good performance for our prompts. To ensure reliability, we attempt to parse the model’s output up to three times. If parsing fails after these attempts, we revert to using the embedding scores as a fallback mechanism. During testing, output-related issues were infrequent; however, they occurred more commonly with larger tables and in scenarios lacking a candidate retriever.
Note that the design of Magneto makes it possible to balance accuracy with computational cost. Based on operational constraints (or requirements), the number of candidates sent to the reranker can be adjusted. The scores for candidates not assessed by the reranker are normalized such that the maximum score aligns with the lowest score received from the reranker. Therefore, a complete ranked list can still be returned to the user.
5. The GDC Benchmark
A key problem in evaluating schema-matching algorithms is the availability of benchmarks that cover a wide variety of challenges present in real data. While there have been efforts to address this issue, benchmarks are often synthetically fabricated from real data (Koutras et al., 2021; Das et al., [n.d.]), which can introduce biases. Moreover, as we show in Section 6 (and as seen in (Koutras et al., 2021)), publicly available benchmarks based on real data such as Magellan (Das et al., [n.d.]) and WikiData (Koutras et al., 2021) are quickly becoming “saturated” since many algorithms attain near-perfect performance and leave small room for algorithmic improvements. This makes it difficult to extract useful insights about the strengths of different algorithms.
To address this problem, we developed a new benchmark dataset based on the real data harmonization scenario described in Example 1.1. We collaborated with experts from a medical school team to design a benchmark that reflects the challenges they face when working with biomedical data. We obtained datasets from ten studies related to tumor analysis (Cao et al., 2021; Clark et al., 2019; Dou et al., 2020; Gillette et al., 2020; McDermott et al., 2020; Huang et al., 2021; Krug et al., 2020; Satpathy et al., 2021; Vasaikar et al., 2019; Wang et al., 2021), and, with the help of the experts, manually aligned and matched these datasets to the Genomics Data Commons (GDC) standard (National Cancer Institute, 2024a).
The GDC is a program of the US National Institutes of Health responsible for receiving, processing, and distributing genomic, clinical, and biospecimen data from cancer research initiatives. Its standard dictionary describes data using a graph model that includes names and descriptions for nodes and attributes, as well as acceptable values for some attributes. Still, for compatibility with existing solutions, we transformed the model to a relational schema that contains only column names and domain information (i.e., we disregard column descriptions). We created a simplified table reflecting the GDC format, the “target” table, listing domain values for each column without repetition. In a future version, it is possible to further extend the benchmark to include multiple tables, and column and value descriptions.
Table 1 illustrates a few samples of the data published by Dou et al. (Dou et al., 2020) alongside their corresponding GDC format. As we can observe, matching clinical data with the GDC standard poses several challenges, including terminology mismatches and data format variations. The benchmark includes 10 pairs of input-target tables. The number of columns in the source tables ranges from 16 to 179, and the number of rows ranges from 93 to 225. Our simplified GDC target schema comprises a single table with 736 columns. While some columns have a small number of distinct values (e.g., binary yes/no attributes), some contain up to 4478 distinct values. The ground truth was manually curated by multiple annotators, who used a mix of manual and automated methods for identifying possible candidate matches (e.g., GDC search tools (National Cancer Institute, 2024b) and bdi-kit (The bdi-kit contributors, 2024)). Given that the correctness of some matches is very challenging to determine even for bioinformatics experts (e.g., it may require reading the original papers or asking data producers), the final match decisions were made by consensus based on what users would expect from an algorithm given the limited context.
6. Experimental Evaluation
6.1. Experimental Setup
Datasets. We evaluated Magneto using the six datasets detailed in Table 3, which we group into two categories: (1) real – human-curated and capture column matches seen in the real world, and (2) synthetic –systematically generated to capture structural variations and diverse match types. The GDC Benchmark (Section 5) reflects challenges in biomedical data integration scenario. The other five datasets are from the Valentine schema matching benchmark (Koutras et al., 2021) and have been used to evaluate recent schema matching solutions (Tu et al., 2023; Du et al., 2024).
Table 3. Statistics of the datasets used for experiments.
Dataset #Table Pairs #Cols #Rows Type
GDC 10 16–733 93–4.5k Real
Magellan 7 3–7 0.9k–131k Real
WikiData 4 13–20 5.4k–10.8k Real
Open Data 180 26–51 11.6k–23k Synthetic
ChEMBL 180 12–23 7.5k–15k Synthetic
TPC-DI 180 11–22 7.5k–15k Synthetic
Baselines. We compare Magneto against several approaches, ranging from traditional string similarity-based methods to sophisticated model-driven techniques. For traditional approaches, we use:
  • $\bullet$
    COMA: Combines multiple strategies to compute and aggregate the similarity of table metadata (Do and Rahm, 2002);
  • $\bullet$
    COMA++: An extension of COMA that leverages column values and their distributions (Aumueller et al., 2005);
  • $\bullet$
    Distribution: Detects column correspondences based on data value distribution  (Zhang et al., 2011);
  • $\bullet$
    SimFlooding: Uses several graph-based techniques to identify correspondences between metadata (Melnik et al., 2002).
We evaluated the Jaccard-Levenshtein baseline proposed in (Koutras et al., 2021) and the Cupid algorithm (Madhavan et al., 2001). However, since they consistently performed significantly worse than the other methods, we exclude them from the reporting for clarity. We used the implementations available in the Valentine repository these methods (Koutras et al., 2021).
We also used recent approaches that use LMs: ISResMat which leverages contrastive learning and embeddings from SLMs (Du et al., 2024), and Unicorn, a supervised and general approach using language models for data encoding and trained for data integration tasks (Tu et al., 2023). We used the implementation from the authors for ISResMat. For Unicorn, we used the pre-trained model and implementation provided by the authors. The match scores were derived from the model’s predicted match probabilities. It is important to note that Unicorn requires external training data for optimal performance. The zero-shot version performed significantly worse than other baselines and was excluded from our evaluation. We report Unicorn numbers for the algorithm in Valentine datasets for completeness, but note that its model was trained on the fabricated datasets from Valentine (Koutras et al., 2021), so data leakage can influence the reported results.
Implementation Details. The COMA algorithm is implemented in Java. Magneto and all other methods are written in Python. We use MPNet as our underlying small language model. It is pre-trained on masked and permuted language tasks, enabling precise contextual understanding of unordered text in column headers and entries (Song et al., 2020; Dong et al., 2023; Sheetrit et al., 2024). For the LLM reranker, we use the GPT-4o-mini model from the OpenAI API due to its robust performance and cost-effectiveness. However, these choices are flexible; alternative models can be easily integrated as replacements based on cost or performance requirements. It is important to note that our study does not focus on determining the optimal model for this task.
We assess four variations of Magneto that represent different combinations of retrievers and rerankers: Magneto-zs-bp (zero-shot SLM, bipartite reranker), Magneto-ft-bp (fine-tuned SLM, bipartite reranker), Magneto-zs-llm (zero-shot SLM, LLM reranker), Magneto-ft-llm (fine-tuned SLM, LLM retriever).
We set margin $m=0.5$ for the fine-tuning configuration, and run the model for 30 epochs on the GDC benchmark and 10 epochs on the Valentine datasets.
Execution platform. The experiments were ran on a server running Red Hat Enterprise Linux version 9.2. We used a single node consisting of 8 cores and 64 GB of memory for CPU-based tasks. The GPU-intensive experiments used an NVIDIA A100 GPU. For Magneto variations, Unicorn, and IsResMat, we used a GPU to run all experiments, except the scalability experiment described in Section 6.3, conducted on a CPU for consistency with other methods.
Figure 6. Magneto attains higher accuracy on the challenging GDC benchmark than traditional, model-based, and supervised approaches to SM.
6.2. End-to-End Accuracy
Performance on GDC. Figure 6 shows the accuracy (measured by MRR and Recall@GT) for all versions of Magneto and the baseline methods for the GDC benchmark. The Magneto variations outperform all other methods for both measures. Magneto-ft-llm has the highest overall accuracy, confirming that the combination of a fine-tuned SLM and an LLM is effective for this complex matching task. The fact that the zero-shot Magneto variations also outperform recent state-of-the-art approaches demonstrates the effectiveness of the serialization and sampling techniques we designed for schema matching. We ablate the different components of Magneto in Section 6.4, where we discuss this in more detail.
Performance on Valentine. The accuracy results for the Valentine benchmark are shown in Figure 7. Note that the different methods attain much higher accuracies for these datasets than for GDC. The Valentine datasets are derived either from open data sources or synthetic repositories, and they are also meticulously curated. In contrast, the GDC benchmark includes real scientific datasets and features a diverse set of column names and values, which, as these results confirm, is particularly challenging for SM solutions.
Figure 7. Accuracy on Valentine datasets. Magneto variants attain high accuracy (even without fine-tuning or data-specific training) compared to traditional, model-based, and supervised approaches.
For all Valentine datasets, Magneto performs on par with or better than the baselines. The only exception is Unicorn. However, as previously noted, this can likely be attributed to data leakage: Unicorn is trained using Valentine datasets. This advantage is particularly apparent in datasets like OpenData: Unicorn achieves significantly higher recall than competing methods. Conversely, for the Magellan dataset, which was not used to train Unicorn, Unicorn has lower accuracy than the other methods, including methods that use only basic metadata, which achieve perfect scores.
The high accuracy of the zero-shot configurations of Magneto leaves little room for improvement. Specifically, fine-tuning faces challenges within the Valentine datasets, which predominantly contains lexical rather than semantic matches. For instance, the OpenData dataset includes matches like Gender to Ge and Employer to Em, which are uncommon in practical applications. Despite these obstacles, our strategy of fine-tuning combined with LLM-based re-ranking generally enhances performance. Additionally, we demonstrate in Section 5 that combining simple synthetic data generation techniques with LLM-derived data leads to better performance in this context.
6.3. Scalability Assessment
We compare the runtime of Magneto to those of the top three baselines in accuracy: Unicorn, Coma++, and ISResMat. Since Coma++ does not support GPU execution, we ran this experiment in CPU mode to ensure a fair comparison. We report only the resuls of Magneto-zs-bp and Magneto-zs-llm since their runtime is similar to Magneto-ft-bp and Magneto-ft-llm, respectively.
For this experiment, we focus on datasets featuring tables with a large number of columns or rows, and we select one source-target pair each from the GDC and OpenData. For the GDC dataset, the source table comprises 179 rows and 153 columns, while the target table contains 4.5k rows and 733 columns. For OpenData, both the source and the target table contain 23k rows and 43 columns. We maintain the input table static and incrementally increase the number of target columns using a random selection. Each execution is repeated 10 times per column number to accommodate randomness. We used a time limit of 2 hours per execution and canceled the executions of any method that exceeded this limit. The results are shown in Figure 8.
Figure 8. Runtime analysis using a log scale on the y-axis. Magneto shows good scalability in scenarios that involve a large number of columns and rows. Magneto-zs-bp is often significantly faster than the other baselines, while Magneto-zs-llm maintains a consistent runtime, completing tasks within minutes, even for datasets with hundreds of columns. Note that ISResMat and Unicorn fail to complete execution within the time limit for GDC after 100 columns.
The performance of Magneto-zs-bp and Magneto-zs-bp is stable as the table size increases; with Magneto-zs-bp showing minimal runtime growth and Magneto-zs-llm demonstrating runtime stability despite its complexity. For GDC, Magneto-zs-bp shows runtimes ranging from 11 to 33 seconds. Such a low increase in runtime reflects its efficient design, bounded primarily by the efficient embedding computations. As expected, Magneto-zs-llm incurs higher runtimes due to the LLM requests. Still, it shows stability for increasing the number of columns with a runtime starting at 545 and peaking at 589 seconds since the amount of data sent does not change as we increase the number of columns in the target. The runtimes of Magneto variations are even lower for OpenData: the dataset has fewer columns but many more rows, which Magneto compensates through sampling.
Coma++ shows low runtimes for a small number of columns, but its performance decreases as the number of columns grows. IsResMat and Unicorn exhibit significant scalability challenges, with their runtime growing substantially with the number of columns. IsResMat was not able to complete the execution for GDC, not even for the lowest number of columns, and its runtime for OpenData was orders of magnitude higher than the other methods. Unicorn could only process the initial 100 columns for GDC.
6.4. Ablation of Magneto Components
Column Serialization. Figure 9 shows the impact of the serialization methods (Section 3.1) on schema matching performance across all datasets based on MRR. We used the serialization strategies on fine-tuned (FT) and zero-shot (ZS) settings: Default, using $\mathcal{S}_{\texttt{default}}$, which combines column names with a small sample of values; Verbose, which uses $\mathcal{S}_{\texttt{verbose}}$ and enriches $\mathcal{S}_{\texttt{default}}$ by incorporating additional instructions; Repeat, which emphasizes header names with $\mathcal{S}_{\texttt{repeat}}$. We also include Header Only-ZS, which uses only column headers without values, serving as a baseline.
In zero-shot settings, $\mathcal{S}_{\texttt{repeat}}$ consistently outperforms other methods, demonstrating the effectiveness of emphasizing column header when training data is absent. $\mathcal{S}_{\texttt{verbose}}$ is more effective in fine-tuned scenarios, demonstrating the benefits of integrating semantic details when domain-specific training is applied. The Header Only approach shows the lowest performance in all datasets, confirming that column headers alone are insufficient for effective schema matching in complex scenarios.
The results further support our assumption that that fine-tuning can greatly improve the performance gains on the GDC dataset. This improvement is likely related to the requirement for external knowledge, which we capture with the LLM-derived training data.
Figure 9. Ablation of column serialization strategies.
Value Sampling. We compare the effectiveness our proposed Priority Sampling (Section 3.1) against the following sampling methods: Coordinated, a variation of Priority that excludes frequency weights; Weighted and Frequency, variations of Prioriry without coordination but with value frequencies – Frequency uses the most frequent values and Weighted uses weighted sampling based on these frequencies; Random uses basic random sampling.
We used GDC and the three synthetic datasets from Valentine; we exclude Wikidata and Magellan since as all methods already perform well on them. A the results in Table 4 show, Magneto-zs-bp with $\mathcal{S}_{\texttt{repeat}}$, the best zero-shot setting configuration, Priority, generally outperforms other methods in both Recall@GT and MRR, and sometimes is a close second. Priority Sampling prioritizes frequently occurring values and enhances the likelihood of sampling similar values across columns, a beneficial feature for schema matching. The top three techniques—including Priority and its ablations—perform comparably well, also representing viable choices.
Synthetic Data Generation. We evaluate our LLM-powered synthetic data generation for SLM fine-tuning against two other methods: simple augmentation, which incorporates row shuffling, sampling, and column name perturbation, and mixed, which combines these methods. We used the optimal settings from serialization studies, $\mathcal{S}_{\texttt{repeat}}$, on Magneto-ft-bp and Magneto-ft-llm. We assess MRR for both methods and Recall@GT for Magneto-ft-bp, as Recall@GT for Magneto-ft-llm is largely influenced by the reranker instead of the fine-tuned candidate retriever. This analysis focuses on the four datasets that have the potential for improvement.
The results in Table 5 indicate that LLM-based and mixed data generation significantly outperform simple perturbations on the challenging GDC dataset. However, on the fabricated Valentine datasets, synthetic training data generated by LLMs shows limited impact due to the datasets nature, which feature uncommon and semantically sparse matches less amenable to LLM-derived enhancements. Nevertheless, the marked performance differential on the GDC dataset highlights the effectiveness of our fine-tuning strategy for real data.
Table 4. Ablation of sampling techniques using Magneto-zs-bp with $\mathcal{S}_{\texttt{repeat}}$. Priority generally outperforms other techniques in Recall@GT and MRR metrics. The best-performing techniques are highlighted in dark blue     , with the second best in light blue    .
Sampling Method GDC ChEMBL OpenData TPC-DI
Recall@GT
Priority 0.344±0.081 0.620±0.264 0.543±0.294 0.726±0.174
Coordinated 0.336±0.069 0.601±0.260 0.506±0.292 0.675±0.224
Weighted 0.342±0.070 0.603±0.266 0.497±0.291 0.643±0.210
Frequency 0.332±0.062 0.525±0.296 0.526±0.300 0.692±0.196
Random 0.334±0.075 0.572±0.268 0.489±0.282 0.665±0.197
MRR
Priority 0.591±0.094 0.900±0.103 0.847±0.200 0.948±0.082
Coordinated 0.586±0.106 0.902±0.105 0.837±0.196 0.937±0.087
Weighted 0.599±0.099 0.888±0.104 0.823±0.212 0.930±0.091
Frequency 0.577±0.104 0.851±0.163 0.833±0.202 0.886±0.130
Random 0.579±0.096 0.885±0.106 0.830±0.195 0.918±0.101
Table 5. Ablation of data generation techniques. These results show that LLM-powered data generation is effective, and better performance can be achieved by combining multiple techniques.
Data Generation GDC ChEMBL OpenData TPC-DI
Recall@GT for Magneto-ft-bp
LLM-based 0.414±0.106 0.785±0.263 0.729±0.271 0.740±0.298
Mixed 0.438±0.085 0.774±0.273 0.743±0.255 0.763±0.255
Simple 0.418±0.099 0.764±0.282 0.711±0.271 0.773±0.261
MRR for Magneto-ft-bp
LLM-based 0.754±0.093 0.932±0.103 0.851±0.152 0.917±0.100
Mixed 0.761±0.071 0.931±0.101 0.841±0.162 0.885±0.125
Simple 0.731±0.099 0.927±0.103 0.817±0.181 0.905±0.117
MRR for Magneto-ft-llm
LLM-based 0.866±0.083 0.960±0.089 0.939±0.080 0.971±0.079
Mixed 0.830±0.077 0.949±0.109 0.927±0.096 0.965±0.096
Simple 0.798±0.116 0.955±0.097 0.917±0.105 0.969±0.081
Match Reranker We performed two experiments to evaluate the match reranker component. The first examines the impact of using the bipartite strategy (Section 2.2). Figure 10 shows the absolute increase/decrease in accuracy from Magneto-zs, which directly returns the retriever matches, to Magneto-zs-bp, which applies the bipartite reranking to the matches produced by the retriever. Note that the bipartite approach can significantly improve Recall@sGT. All datasets benefit from the technique at some point, except Magellan, whose scores are already (very) high. The improvements can reach nearly 0.6 in absolute terms. The mean improvement varies across datasets; average recall scores across datasets other than Magellan range from 0.03 (GDC) to 0.17 (OpenData). The improvements for MRR are lower but still sizeable for ChemL and Open Data. While reranking with bipartite generally works, we observed that it can occasionally lead to decreased MRR and Recall@GT.
Our final experiment evaluates: (1) how the LLM-based reranker improves upon the traditional bipartite approach, and (2) how its performance varies as we increase the number of match candidates sent to the LLM. We focus on the GDC dataset, which has a higher column count. For this experiment, we used the zero-shot settings for the retriever. Figure 11 shows accuracy and runtime for different numbers of candidates ($k$).
Figure 10. Ablation of rerankers. Absolute accuracy improvement using Magneto-zs-bp over Magneto-zs.
Figure 11. Ablation of number of candidates sent to LLM-based reranker on GDC.
The LLM-based re-ranking approach improves over the bipartite. For example, at $k=5$, we observe a 12.3% absolute improvement in Recall@GT over the bipartite baseline, rising from 0.404 to 0.454. Mean Reciprocal Rank (MRR) demonstrates even more promising gains, from 0.551 to 0.605—a 9.8% increase. While MRR continues to improve with more columns, processing all target table columns becomes computationally infeasible for large schemas. That is why we argue for combining a retriever with an LLM-based re-ranker for schema matching. Note that candidates that are not forwarded to the reranker are assigned lower scores, this leads to a decrease in Recall@GT since it considers all pairs based on these scores.
As the number of columns increases, computational costs grow. The runtime reflects the latency of API-based LLM requests. However, we can expect similar trends with alternative LLM approaches, as prompts with increasing length become more challenging for LLMs (Wan et al., 2024). Still, the runtime remains practical and competitive for usable values of $k$, as demonstrated in our scalability experiment. We used $k=20$ for our other experiments, as it balances good MRR, Recall@GT, and practical runtime across GDC and other datasets.
7. Related Work
A wide range of methods have been proposed for schema matching, an important research area in the database field (Rahm and Bernstein, 2001). Related approaches to our work can be broadly categorized into three main groups: traditional methods, small language model-driven methods, and large language model-based methods. Below, we review representative work in each category.
Traditional Methods. Over the years, researchers have developed and combined various approaches for automatic schema matching (Bellahsene et al., 2011; Doan et al., 2012; Koutras et al., 2021). A straightforward yet effective approach is detecting overlap in column names (Madhavan et al., 2001; Melnik et al., 2002; Do and Rahm, 2002) and overlap in column values (Aumueller et al., 2005; Zhang et al., 2011). Some incorporate relaxations when measuring overlaps, such as accounting for syntactic and semantic similarities between column names and values (Madhavan et al., 2001; Aumueller et al., 2005). Others also consider factors like data type relevance and value distribution (Aumueller et al., 2005; Zhang et al., 2011). Among these, the COMA algorithm stands out for integrating the most strategies and weighting their outputs to achieve better accuracy (Do and Rahm, 2002; Aumueller et al., 2005), often remaining competitive even against more recent approaches (Koutras et al., 2021).
However, these approaches often struggle to capture complex relationships and deeper semantics within datasets, resulting in poor accuracy when identifying semantically valid matches (Khatiwada et al., 2023).
Small Language Model-Driven Methods. Small Language Models (SLMs) have transformed schema matching by using embeddings similarity (Cappuzzo et al., 2020), effectively encoding and comparing textual and contextual data within columns. Furthermore, contrastive learning refines SLMs’ ability to distinguish between matching and non-matching column pairs (Fan et al., 2023; Cong et al., 2023; Zhang et al., 2023). Meanwhile, synthetic tabular data generation facilitates self-supervised learning, allowing models to improve without ground truth (Fan et al., 2023; Du et al., 2024).
Among these methods, ISResMat customizes pre-trained models for dataset-specific adaptation, generating training pairs from table fragments, and applying pairwise comparison losses to refine matching accuracy (Du et al., 2024). It also addresses cardinality constraints and manages numerical columns with advanced techniques; Unicorn, a state-of-the-art general matching model, employs contrastive learning and a Mixture-of-Experts (MoE) layer within its architecture to discern matches but relies on supervised training (Tu et al., 2023).
Large Language Model-Driven Methods. Recent works have leveraged large language models (LLMs) for various aspects of tabular data management, predominantly focusing on single-table tasks (Kayali et al., 2024; Li et al., 2024a; Feuer et al., 2024; Hegselmann et al., 2023). Some studies discuss LLM applications in schema matching, highlighting the potential of LLMs in this area. However, these approaches often rely solely on prompting strategies—either fine-tuned or zero-shot—which suffer from scalability issues and high computational costs (Li et al., 2024a; Parciak et al., 2024; Fang et al., 2024; Feng et al., 2024).
Furthermore, Sheetrit et al. (2024) and Zhang et al. (2024) utilize zero-shot pre-trained LLMs for schema matching. Sheetrit et al. (2024) addresses multiple-table matching, whereas our work focuses on the less-explored two-table schema matching, particularly for tables with numerous columns. Zhang et al. (2024) uses rule-based feature extraction and trains an XGBoost classifier with gold data, a supervised approach distinct from our unsupervised method. Hence, we do not compare our findings directly with these studies.
8. Conclusions
We have proposed Magneto, a novel framework that leverages small and large language models to derive schema matching strategies that generalize across domains and balance accuracy and runtime tradeoffs. We introduce a new benchmark that captures some of the complexities involved in biomedical data integration and that also presents new challenges for schema matching methods. Through a detailed experimental evaluation, including comparisons against state of the art methods and ablations, we demonstrate the effectiveness of Magneto and our design choices.
The source code for all baselines and Magneto can be accessed from https://github.com/VIDA-NYU/data-integration-eval and https://github.com/VIDA-NYU/data-harmonization-benchmark.
Acknowledgments.
We thank our collaborators from the NYU School of Medicine, David Fenyo, Wenke Liu, and Sarah Keegan, for sharing their experiences in biomedical data integration and contributing to the creation of the GDC benchmark. We would like to thank DARPA for funding this work through the Automating Scientific Knowledge Extraction and Modeling (ASKEM) program, Agreement No. HR0011262087. This work was also partially supported by NSF award IIS-2106888. The views, opinions, and findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the Department of Defense, the U.S. Government, or NSF. We also thank Roque Lopez, Benjamin Feuer, and Grace Fan for their constructive suggestions.